OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. (5 points) Part B: The length of a rectangle is 7 units and its width is square root of 7 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Ganpat

OpenStudy (anonymous):

I know Part A is... @Ganpat

OpenStudy (ganpat):

Area of rectangle = l * w A = 7 * (root) 7 = 7 So its like \[7 \sqrt{7}\]

OpenStudy (ganpat):

probably it is irrational number, it doesn't have a specific ratio.. I am not very much sure...

OpenStudy (anonymous):

the first one you post, is that for part b?

OpenStudy (ganpat):

yeah tha'ts a part B..

OpenStudy (anonymous):

here's what I got on part A, Where a number is raised to a fraction it can be converted to a radical. The 2 in 2√64 would become the denominator and a 1 would be the numerator changing it to 64.

OpenStudy (anonymous):

Thanks @Ganpat